### v1

# FedSDAF: Leveraging Source Domain Awareness for Enhanced Federated Domain Generalization

Abstract.
Traditional domain generalization approaches predominantly focus on leveraging target domain-aware features while overlooking the critical role of source domain-specific characteristics, particularly in federated settings with inherent data isolation. To address this gap, we propose the Federated Source Domain Awareness Framework (FedSDAF), the first method to systematically exploit source domain-aware features for enhanced federated domain generalization (FedDG). The FedSDAF framework consists of two synergistic components: the Domain-Invariant Adapter, which preserves critical domain-invariant features, and the Domain-Aware Adapter, which extracts and integrates source domain-specific knowledge using a Multihead Self-Attention mechanism (MHSA). Furthermore, we introduce a bidirectional knowledge distillation mechanism that fosters knowledge sharing among clients while safeguarding privacy. Our approach represents the first systematic exploitation of source domain-aware features, resulting in significant advancements in model generalization capability.Extensive experiments on four standard benchmarks (OfficeHome, PACS, VLCS, and DomainNet) show that our method consistently surpasses state-of-the-art federated domain generalization approaches, with accuracy gains of 5.2-13.8%. The source code is available at https://github.com/pizzareapers/FedSDAF.
Federated learning, Cross-domain, Federated Domain Generalization, Domain-aware features
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 03–05, 2025; Dublin, Irelandsubmissionid: 5352ccs: Computing methodologies Machine learningccs: Computing methodologies Distributed computing methodologies
Figure 1. Ilustration of Domain Generalization.
[…]
A woman and a girl in white dresses sit in an open car.
1. Introduction
The synergy between federated learning (FL) and large-scale pre-trained models(Houlsby et al., 2019; Su et al., 2024b) has ushered in a new era of distributed learning paradigms for privacy-sensitive applications. While foundational architectures, particularly Transformers, demonstrate remarkable cross-task adaptability(Mahla et al., 2025; Chen et al., 2022) through parameter efficient techniques, conventional FL frameworks(McMahan et al., 2016; Wang et al., 2020) encounter a fundamental limitation: their underlying independent and identically distributed (i.i.d.) assumption stands in stark contrast to the inherent multi-source domain heterogeneity in real-world scenarios(Sahu et al., 2018; Peng et al., 2020; Liu et al., 2021a). This misalignment presents a dual challenge: not only do client-specific data emerge from distinct domain distributions, but models must also simultaneously address inter-client domain shifts while maintaining robust generalization capabilities for unseen target domains during deployment. These intertwined challenges constitute the core problem in Federated Domain Generalization (FedDG).
Standard domain generalization (DG) methods have advanced through Domain-invariant representation learning(Tan et al., 2024; Peng and Chan, 2024) and Feature disentanglement(Cheng et al., 2024). Both Feng et al(Feng et al., 2024a). and Li et al(Li et al., 2024b). have adopted the Feature Disentanglement approach, which separates features into domain-invariant and domain-specific components, utilizing domain-aware features from the target domain to enhance the model’s generalization ability in text classification and target detection tasks. However, these studies primarily focus on leveraging domain-aware features in the target domain, without fully exploring the potential value of domain-aware features in the source domain.
Federated Domain Generalization (FedDG) extends beyond the conventional challenges of standard domain generalization by introducing a critical constraint: the prohibition of raw data sharing across clients. This restriction impedes explicit multi-domain alignment through centralized techniques such as feature matching or adversarial training, thereby limiting the models’ ability to fully leverage knowledge from diverse source domains. Current FedDG approaches have attempted various solutions: Liu et al.(Liu et al., 2021b) facilitated domain knowledge sharing through the exchange of amplitude spectrum information in frequency space; Huang et al.(Huang et al., 2023) implemented class prototype sharing derived from local samples; and Chen et al.(Chen et al., 2023) enabled cross-client style transfer through the exchange of pixel-level feature channel statistics. While these methods represent progress in addressing data isolation, they primarily focus on transferring fragmentary domain characteristics derived directly from local sample statistics, leaving the potential of source domain-aware features largely unexplored.
In the context of federated learning, data isolation fundamentally constrains the model’s ability to effectively utilize knowledge from diverse domains. While existing methods transmit fragmented domain characteristics (e.g., spectral signatures or prototypes), these incomplete knowledge representations fail to capture the crucial relationships between domain-specific patterns and underlying task semantics, resulting in suboptimal cross-domain knowledge integration during model aggregation. Notably, our controlled experiments in Section 3.2 reveal a counterintuitive finding: source domain-aware features demonstrate superior generalization capabilities compared to target domain-derived features in data-isolated scenarios. This observation suggests that the predominant focus of existing methods on target domain adaptation may overlook the valuable contribution of source domain-specific patterns in developing transferable representations.
To address these limitations, we introduce the Federated Source Domain Awareness Framework (FedSDAF), which systematically leverages source domain-aware features through dual adaptive components. The first component, a Domain-Invariant Adapter, preserves essential cross-domain patterns, while the second component, a Domain-Aware Adapter, employs Multi-head Self-Attention (MHSA) to dynamically integrate source-specific knowledge with base model representations, enabling hierarchical knowledge integration. Furthermore, the framework incorporates a bidirectional knowledge distillation mechanism that facilitates logarithmic knowledge exchange between clients and the server, thereby enabling effective knowledge transfer from diverse source domains to the global model. This innovation alleviates the challenges posed by data isolation constraints. Our framework not only enhances the model’s comprehension of domain-specific patterns but also improves cross-domain knowledge integration, ultimately boosting the model’s ability to generalize across multiple target domains. Our contributions are summarized below:
  • We introduce a novel framework, FedSDAF, which systematically addresses the inherent limitations of data isolation in Federated Domain Generalization by leveraging source domain-aware features. The framework consists of dual adaptive components—Domain-Invariant Adapter and Domain-Aware Adapter—that work in tandem to preserve essential cross-domain patterns while dynamically integrating source-specific knowledge. This approach enhances the model’s understanding of domain-specific intricacies, fostering improved generalization capabilities across multiple target domains.
  • We introduce a bidirectional knowledge distillation mechanism that enables knowledge exchange between clients and the server, effectively alleviating the data isolation constraints in federated learning while facilitating comprehensive knowledge transfer from diverse source domains to the global model.Development of a Novel Framework for Leveraging Source Domain-Aware Features.
  • We demonstrate the framework’s effectiveness through extensive experiments on four benchmark datasets, achieving state-of-the-art performance in domain generalization tasks.
2. Related Work
2.1. Federated Learning
Federated Learning represents a paradigm shift in distributed model optimization, enabling collaborative training across multiple devices or organizations without sharing raw data(Tang and Wang, 2023; Fu et al., 2024). The fundamental principle involves local model training on client devices followed by parameter aggregation at a central server, creating a global model that benefits from diverse data sources while maintaining data privacy(Gad et al., 2024; Cai et al., 2024). One of the primary challenges in federated learning environments is the non-Independent and Identically Distributed (non-IID) nature of client data, which significantly impacts model performance(Islam et al., 2024; Feng et al., 2024b). To address these challenges, researchers have proposed various schemes and models. For instance, Zhang et al(Zhang et al., 2022). introduced a method for personalizing local models, while Guo et al(Guo and Yang, 2024). developed a more robust algorithm aimed at learning more efficient global models(Li et al., 2024a). Despite these advances, most research on non-IID challenges focuses on addressing out-of-sample gaps in unseen client data with known distributions. However, as Yuan et al. (Yuan et al., 2021) highlight, there is still a critical gap in handling unseen client distributions(Jiang et al., 2024). This highlights the need for approaches that can generalize beyond the training distributions, especially when encountering entirely new domains.
2.2. Domain Generalization
The primary objective of domain generalization is to enable models to perform well on unseen domains by learning robust representations from multiple source domains. Within this framework, representation learning has emerged as a pivotal approach, concentrating on the extraction of domain-agnostic features that generalize across different distributions(Lv et al., 2022; Liu et al., 2024a). Existing research in representation learning for domain generalization can be broadly categorized into two main strands: Domain-Invariant Representation Learning (DIRL)(Xu et al., 2022) and Feature Disentanglement(Gholami et al., 2023). DIRL methods aim to extract features that remain consistent across domains through techniques such as kernel tricks, adversarial learning, and explicit feature alignment. These approaches strive to minimize the distribution discrepancy between source domains, operating under the assumption that domain-invariant representation will transfer effectively to unseen target domains(Su et al., 2024a). In contrast, feature disentanglement involves dividing input features into domain-invariant and domain-specific components. This separation allows models to isolate generalizable patterns from domain-specific variations, potentially enhancing performance on unseen domains. Significant contributions in this area include the work of Feng et al(Feng et al., 2024a). and Li et al(Li et al., 2024b)., who employed feature disentanglement to improve model generalization in text classification and target detection tasks. Additionally, they used features specific to the target domain, combined with domain-invariant features, to improve the model’s performance on that domain. However, a notable limitation of current feature disentanglement approaches is their predominant focus on utilizing domain-aware features from the target domain, while not fully exploring the potential value of domain-aware features from source domains(Zhang et al., 2023a). This represents a significant opportunity, particularly in federated scenarios where data isolation presents additional challenges.
2.3. Federated Domain Generalization
Federated Domain Generalization (FedDG) is an emerging research area that addresses the challenges posed by traditional federated learning (FL) and domain generalization (DG) approaches(Zhang et al., 2023d). The primary objective of FedDG is to tackle data isolation in federated scenarios while enhancing the generalization capabilities of models across unseen target domains(Radwan and Shehata, 2024). For instance, Liu et al(Liu et al., 2021b). proposed a method that leverages the exchange of frequency-space magnitude spectral information. Similarly, Wan et al(Wan et al., 2024). facilitated the sharing of class prototypes extracted from local samples among different clients. Furthermore, Chen et al. enabled cross-client style migration by exchanging the mean and variance of pixel-level feature channels. Recent studies, such as FedGCA(Liu et al., 2024b), have also introduced global consistency augmentation to address the challenges associated with single-source federation domain generalization. Although these approaches mitigate the issue of data segregation to some extent, they still exhibit significant limitations, The exchanged domain knowledge components (spectral signatures, prototypes, style parameters) operate at shallow feature levels, failing to capture deep semantic correlations between domain-specific patterns and task objectives.
3. Methodology
3.1. Problem Setup
Traditional Domain Generalization (DG): Consider $K$ source domains ${\mathcal{S}^{k}}{k=1}^{K}$ with distinct distributions $P^{k}\neq P^{k^{\prime}}\ (k\neq k^{\prime})$, where each domain consists of labeled samples $\mathcal{S}^{k}={(x_{i}^{k},y_{i}^{k})}{i=1}^{n_{k}}$. The objective is to learn a model $f_{\theta}:\mathcal{X}\to\mathcal{Y}$ that minimizes the expected risk on an unseen target domain $\mathcal{T}$ with distribution $P^{\text{tgt}}$:
(1) $\min_{\theta}\mathbb{E}{(x,y)\sim P^{\text{tgt}}}\mathcal{L}(f\theta(x),y)$
Federated Domain Generalization (FedDG): Extends the traditional DG framework by incorporating a critical privacy constraint :
Data Isolation: Source domains are distributed across $K$ clients ${\mathcal{C}k}{k=1}^{K}$, where each client’s data must remain private and locally contained.
The FedDG objective can be formulated as a constrained optimization problem:
(2) $\begin{split}\min_{\theta_{g}}&\underbrace{\mathbb{E}_{(x,y)\sim P^{\text{tgt}}}\mathcal{L}(f_{\theta_{g}}(x),y)}_{\text{Target Risk}}\\ &\text{s.t.}\quad\underbrace{\theta_{g}=\mathcal{A}(\theta_{k}),\ \mathcal{S}^{k}\subseteq\mathcal{C}_{k}}_{\text{Federated Constraints}}\end{split}$
3.2. Motivational study
Building upon the problem setup of federated domain generalization , we pose a critical research question: Which type of domain-aware features contributes more significantly to model performance in target domains - those derived from the source domain or the target domain itself? To investigate this, we conduct controlled experiments on the PACS dataset using a leave-one-domain-out protocol. Specifically, we train domain-aware adapters independently on three source domains (Photo, Art, Cartoon) while keeping the domain-invariant adapter fixed. Each adapter is then evaluated on all possible target domains without any target domain fine-tuning. Table 1 reveals a counterintuitive finding: when testing on a target domain (e.g., Sketch), adapters trained on other source domains consistently outperform those trained on the target domain itself. For instance, the Art-trained adapter achieves 97.07% accuracy on Sketch, surpassing the Sketch-trained adapter’s 94.94%.
Target Domain P A C
Adapter in Domain P 96.49% 97.07% 97.82%
Adapter in Domain A 98.86% 89.42% 97.35%
Adapter in Damain C 98.26% 94.38% 94.94%
Table 1. Accuracy results of the domain-aware perceptual adapters combined with the global domain-invariant adapter across different target domains in the PACS dataset.
This phenomenon can be explained through two key perspectives from our introduction: (1) Data isolation constraints in federated learning prevent effective collaborative modeling of target domain features, making their isolated extraction suboptimal; (2) Source domain features provide complementary contextual knowledge that helps bridge domain gaps through their inherent diversity. These findings motivate our core hypothesis: Systematic exploitation of source domain-aware features through collaborative federation can create more robust representations for unseen target domains. This insight drives our framework design that strategically aggregates source domain knowledge while preserving domain-specific characteristics through novel adaptation mechanisms.
3.3. Proposed Method
In this section, we formally introduce our proposed Federated Source Domain-Aware Framework (FedSDAF). The approach comprises two key components: a Knowledge Integrator Adapter (KIA) and a Bidirectional Knowledge Distillation (BKD) process.
Figure 2. FedSDAF Architecture Design.
[…]
A woman and a girl in white dresses sit in an open car.
3.3.1. Knowledge Integrator Adapter (KIA)
The core idea of KIA is to explicitly disentangle and leverage two complementary representation components: domain-invariant features shared by all clients and domain-specific features unique to each client. To achieve this, we introduce two types of adapters into a shared backbone neural model: a Domain-Invariant Adapter $A_{di}$ and a Domain-Aware Adapter $A_{dw}$.
(1) Domain-Invariant Adapter:The Domain-Invariant Adapter $A_{di}$ aims to capture stable and generalizable representations across all source domains. At each federated communication round $r$, the server maintains a global domain-invariant adapter $A_{di}^{r}$, which is distributed to all clients for local fine-tuning. Each client $k$ optimizes a local copy of this adapter, denoted as $\hat{A}_{di}^{k,r}$, on its private dataset. After local training, the server aggregates these adapters into an updated global adapter using a weighted averaging scheme, which is essential for maintaining a robust representation that can effectively generalize across various domains.
(2) Domain-Aware Adapter: Before the first round of communication, each client initializes a domain-aware adapter $A_{dw}$ locally. Unlike the domain-invariant adapter, this domain-aware adapter is updated strictly locally throughout the training process to force it to capture the domain-aware features of each domain.. Its role is to capture subtle, client-specific knowledge from the local domain, thus preserving rich domain-specific information. The domain-aware adapter is crucial for enhancing the model’s ability to adapt to the specific characteristics of the client’s data.
The Knowledge Integrator Adapter (KIA), which integrates both the Domain-Invariant Adapter and the Domain-Aware Adapter (as illustrated in Fig. 2), is inserted following the Feedforward Neural Networks (FFNs) layer of the neural network. To effectively integrate these two sets of features, we employ a Multi-Head Self-Attention (MHSA) mechanism. Let $h$ denote the intermediate representation output from the backbone model. The integrated feature representation on client $k$ is computed as follows:
(3) $\hat{A}_{dw}\left(h\right)=MHSA\left(LN\left[A_{dw}\left(h\right);h\right]\right)$
(4) $h^{\prime}=h+\frac{1}{2}\hat{A}_{dw}(h)+\frac{1}{2}\hat{A}_{di}(h)$
This integration ensures that in each round of communication, the model effectively assimilates knowledge from the source domain while leveraging the rich contextual information inherent in the domain-aware features. By utilizing MHSA, the model can dynamically weigh the importance of various features from both the source domain and the base model’s outputs, facilitating the formation of more adaptive domain-aware features.
Figure 3. Optimization process
3.3.2. Bidirectional Knowledge Distillation (BKD)
The bidirectional knowledge distillation process serves as the core mechanism within the FedSDAF framework, designed to facilitate effective knowledge sharing among various clients. This process enables the model to transfer and integrate knowledge across different source domains, thereby achieving a comprehensive utilization of knowledge from multiple sources. Specifically, the bidirectional knowledge distillation involves the following steps: First, two Kullback-Leibler (KL) divergences are computed using the following formulas:
(5) $L_{di}^{KL}=KL(z_{di}\circ(x)\parallel\hat{z}_{dw}\circ(x))$
(6) $L_{dw}^{KL}=KL(\hat{z}_{dw}\circ(x)\parallel z_{di}\circ(x))$
Here, KL denotes the Kullback-Leibler divergence, while $z_{di}$ and $\hat{z}_{dw}$ represent the prediction logits of the base model injected with $A_{di}$ and $\hat{A}_{dw}$, respectively. Through this setup, bidirectional knowledge distillation allows clients to exchange knowledge, facilitating learning and adaptation across different source domains. By sharing prediction logits, the model is able to acquire additional domain information from the server pertaining to other domains. This interaction of information is crucial in compensating for knowledge gaps arising from data isolation.
During the actual training process, we employ a distillation loss in conjunction with Cross-Entropy Loss (CEL), introducing a weighting factor to maintain a balance between the two losses,Figure 3 illustrates the process of bi-directional distillation:
(7) $L=L^{CE}+\alpha L^{KL}$
3.3.3. Optimization process
As depicted in Algorithm 1, the updating process occurs in two phases: server-side and client-side. On the server side, the global domain-invariant adapter $A_{di}^{0}$ and the domain-aware adapters $A_{dw}^{k,0}$ for each client are initialized along with the initial weights $a_{k}^{0}$. In each communication round, if it is not the initial round, the server collects the generalization gap $G_{k}^{r}$ and updated domain-invariant adapters $A_{di}^{k,r}$ from all clients. The server then adjusts the weights $a_{k}^{r}$ based on client performance and aggregates this information using the GA strategy(Zhang et al., 2023c), represented as:
(8) $A_{\mathrm{inv}}^{(r+1)}=\sum_{k=1}^{K}\frac{G_{k}^{(r)}}{\sum_{j=1}^{K}G_{j}^{(r)}}\hat{A}_{\mathrm{inv},k}^{(r)},$
This iterative process of adapter refinement and knowledge distillation culminates in the broadcast of the updated global domain-invariant adapter to all clients, completing a synergistic cycle of domain-invariant feature consolidation and source-specific knowledge preservation. The continuous interplay between these dual adapters across communication rounds progressively enhances the model’s capacity to disentangle cross-domain patterns while retaining discriminative source characteristics.
4. Experimental Results
In this section, we describe the details of our experimental setup and report on a series of experimental results to illustrate the validity of our approach.
4.1. Datsets
We evaluate the effectiveness of our method using four mainstream domain generalization datasets. Below is a description of each dataset.
PACS Dataset(Li et al., 2017): The PACS dataset consists of 9,991 images across seven object categories distributed among four visually distinct domains: Photo, Art Painting, Cartoon, and Sketch.
OfficeHome Dataset(Venkateswara et al., 2017): The OfficeHome dataset comprises 24,788 images spanning 65 categories across four domains: Art, Clipart, Product, and Real World.
VLCS Dataset(Fang et al., 2013): The VLCS dataset integrates images from four photographic collections: PASCAL VOC2007, LabelMe, Caltech-101, and SUN09, containing a total of 10,729 images across five common categories. .
DomainNet Dataset(Peng et al., 2019): DomainNet is the largest and most challenging benchmark for domain generalization, comprising 569,010 images across 345 categories distributed over six domains: Clipart, Infograph, Painting, Quickdraw, Real World, and Sketch.
For each of the datasets mentioned above, we employed a leave-one-domain-out testing strategy. Specifically, one domain is selected as the target domain for testing while the remaining domains serve as source domains for model training. Finally, we report the classification accuracy and overall average accuracy for each domain.
4.2. Baseline Methods
We comprehensively compared our FedSDAF approach against three main categories of methods, encompassing a total of 14 baseline approaches:
(1) Centralized learning based Domain Generalization (DG) Methods. These methods operate by centrally collecting data from all available domains to train models with robust generalization capabilities. Specifically, we included the following approaches in this category: SWAD(Cha et al., 2021), which identifies flat minima through systematic random weight averaging, thereby stabilizing the learning process across different domains and enhancing generalization performance; HCVP(Zhou et al., 2025), which employs a hierarchical cue generation system that effectively integrates domain and task features, consequently enhancing cross-domain performance; and Doprompt(Zheng et al., 2022), which generates adaptive cues specifically tailored to target images, thus significantly improving generalization capabilities when applied to previously unseen domains.
(2) CNN-based Federated Domain Generalization (FedDG) Methods. These methods are based on the CNN architecture specifically developed to address the data segregation challenge in joint learning scenarios. We evaluated: FedSR(Nguyen et al., 2022), which learns generalized representations by effectively leveraging L2 regularity and mutual information constraints among federated clients, thereby promoting knowledge sharing while respecting data privacy; FedADG(Zhang et al., 2023b), which utilizes adversarial learning techniques to align source domains with reference distributions in a federated context, consequently enhancing domain generalization; CCST(Chen et al., 2023), which implements cross-client style migration to enable sharing of domain-specific knowledge between clients without directly exchanging raw data; ELCFS(Liu et al., 2021a), which facilitates information exchange between clients through frequency domain transformations, thus preserving privacy while enabling effective knowledge transfer; and GA(Zhang et al., 2023c), which incorporates a dynamic weight adjustment module on the server side to optimize the global model’s generalization capability across diverse domains.
(3) Parameter Efficient Fine-Tuning (PEFT) Methods. These methods integrate adapters for efficient fine-tuning, primarily aimed at improving model generalization in visual-language models (VLMs). Our comparison included: FedCLIP(Lu et al., 2023), which employs a CLIP-based attention adapter that effectively leverages visual-language alignment within a federated context; PromptFL(Guo et al., 2024), which learns client-independent textual cues after aggregation, enabling more efficient knowledge sharing while maintaining model performance; FedAPT(Su et al., 2024c), which collaboratively trains meta-cues and adaptive networks to significantly enhance generalization capabilities across heterogeneous domains; FedPR(Feng et al., 2023), which strategically optimizes local visual cues within a global cue zero space, achieving an optimal balance between local adaptation and global consistency; FedMaPLe(Khattak et al., 2023), which implements multimodal cue learning in a federated environment to effectively leverage diverse information sources; and PLAN(Gong et al., 2024), an innovative framework that utilizes reference-based prompt learning to synchronize local prompts across clients while selectively aggregating them into global prompts, thereby substantially enhancing generalization in federated domain generalization scenarios.
Table 2. Performance Comparison on PACS, VLCS, and OfficeHome Datasets (%)
Methods PACS Avg. VLCS Avg. OfficeHome Avg.
A C P S C L V S A C P R
Centralized learning based Domain Generalization Methods
SWAD $93.23$ $85.93$ $99.18$ $82.03$ $90.44$ $98.49$ $68.36$ $75.40$ $79.49$ $79.31$ $76.26$ $68.87$ $86.74$ $87.03$ $79.73$
HCVP $93.17$ $86.89$ $99.33$ $81.30$ $90.17$ $96.32$ $66.26$ $76.40$ $81.65$ $81.08$ $81.77$ $69.76$ $88.01$ $90.62$ $82.54$
Doprompt $95.00$ $86.35$ $99.63$ $78.20$ $89.91$ $96.70$ $66.53$ $78.28$ $79.39$ $80.23$ $80.95$ $70.88$ $88.94$ $90.10$ $82.72$
CNN-based FedDG Methods(backbone:Resnet-50)
FedSR $88.19$ $67.45$ $95.74$ $65.92$ $79.33$ $95.16$ $65.86$ $78.51$ $73.49$ $78.26$ $69.12$ $49.69$ $72.71$ $79.12$ $67.66$
FedADG $82.93$ $65.42$ $98.09$ $65.36$ $77.95$ $95.21$ $65.76$ $76.43$ $75.96$ $78.34$ $69.31$ $48.76$ $72.89$ $79.13$ $67.52$
CCST $87.02$ $74.57$ $98.29$ $65.84$ $81.43$ $96.49$ $65.73$ $76.42$ $77.67$ $79.08$ $69.23$ $51.36$ $72.09$ $81.27$ $68.19$
ELCFS $86.77$ $73.21$ $98.14$ $65.16$ $80.82$ $95.67$ $65.02$ $76.55$ $77.96$ $79.80$ $68.17$ $50.52$ $71.44$ $80.11$ $67.56$
GA $87.68$ $75.19$ $97.56$ $65.86$ $81.57$ $96.77$ $65.16$ $78.89$ $78.93$ $79.18$ $68.62$ $50.60$ $73.35$ $81.23$ $68.45$
(PEFT) Methods(backbone: ViT-B/16)
FedCLIP $96.19$ $97.91$ $99.76$ $85.85$ $94.93$ $99.93$ $66.98$ $73.28$ $87.14$ $81.83$ $78.45$ $64.77$ $73.28$ $87.84$ $79.69$
PromptFL $96.34$ $98.46$ $99.58$ $92.19$ $96.64$ $99.71$ $68.03$ $72.24$ $85.10$ $83.59$ $82.98$ $68.98$ $92.14$ $90.27$ $83.59$
FedAPT $97.15$ $99.12$ $99.69$ $92.34$ $97.08$ $99.36$ $68.18$ $81.06$ $85.98$ $83.64$ $83.96$ $71.65$ $91.93$ $90.51$ $84.51$
FedPR $98.10$ $99.02$ $99.88$ $91.11$ $97.03$ $99.36$ $68.18$ $81.06$ $85.98$ $83.64$ $84.04$ $71.63$ $92.39$ $91.34$ $84.58$
FedMaPLe $98.44$ $99.02$ $99.94$ $90.40$ $96.95$ $98.02$ $69.54$ $82.15$ $85.81$ $83.87$ $84.56$ $72.82$ $92.38$ $91.07$ $85.21$
PLAN $98.58$ $99.14$ 99.82 $92.08$ $97.40$ 99.18 $69.94$ $83.75$ $88.28$ $85.29$ $86.65$ $74.73$ $93.47$ $92.06$ $86.73$
FedSDAF 98.88 99.36 $99.28$ 96.56 98.52 $98.72$ 85.09 87.38 94.33 91.38 87.15 76.20 94.38 92.78 87.63
Note: Bold numbers indicate the best performance in each category. All values are in percentage (%).
Table 3. Performance Comparison on DomainNet Datasets (%)
Methods DomainNet
C I P Q R S Avg.
FedCLIP $74.12$ $48.36$ $68.49$ $31.73$ $80.52$ $58.62$ $60.31$
PromptFL $76.53$ $51.72$ $70.86$ $34.21$ $81.68$ $68.37$ $63.90$
FedAPT $77.02$ $51.45$ $70.36$ $49.62$ $86.64$ $68.43$ $67.25$
FedPR $75.49$ $51.96$ $71.42$ $35.98$ $82.67$ $69.43$ $64.49$
FedMaPLe $78.61$ $65.23$ $71.89$ $43.46$ $86.32$ $72.46$ $69.67$
PLAN $79.51$ 66.42 $72.11$ $48.83$ 86.72 $72.69$ $71.05$
FedSDAF 82.59 $62.95$ 85.12 54.16 $84.99$ 74.14 73.99
4.3. Implementation Details
In our implementation of FedSDAF, we employed the ViT-B16 architecture as the backbone network. The federated learning process was configured to span 200 communication rounds, with each client performing 30 local epochs per round. To accommodate the diverse characteristics of the benchmark datasets, we utilized dataset-specific batch sizes: 128 for both PACS and Office Home datasets, 64 for VLCS, and 1024 for the more extensive DomainNet. For optimization, we adopted the Scaffold optimizer with an initial learning rate of 0.001 to facilitate effective coordination of client model updates. To regulate the learning rate throughout the training process, we implemented PyTorch’s StepLR scheduler with a step size of 0.2 and a decay factor of 0.1, ensuring appropriate learning rate adjustments during model convergence.
4.4. Main Results
Tables 2 to 3 show the comparison of our proposed FedSDAF method with other methods on PACS, VLCS, Officehome and DomainNet, respectively.
In our experimental evaluations, we achieved state-of-the-art (SOTA) results across all four mainstream datasets: PACS, VLCS, OfficeHome, and DomainNet. This accomplishment underscores the robust generalization capabilities of our proposed Federated Source Domain Awareness Framework (FedSDAF). Notably, our method demonstrated significant advancements on the S-domain of the PACS dataset, where previous approaches had struggled to deliver satisfactory outcomes. Specifically, our framework achieved an impressive accuracy of 98.62%, reflecting a notable improvement of 3.6
Furthermore, we observed strong performance on DomainNet, particularly in the Quickdraw domain, where our method excelled despite the inherent challenges posed by unnatural image features. This success further validates our framework’s adaptability to diverse and complex data distributions. The VLCS dataset, which integrates images from varying photographic styles, highlights the differences in visual representation that our model adeptly addressed. By effectively leveraging domain-aware features from the source domains, our method capitalized on the nuances of each dataset, thereby enhancing its performance across the board.
Standard domain generalization (DG) methods, while demonstrating excellent performance across various datasets, rely on centralized frameworks that necessitate unrestricted access to client data, which compromises data privacy. This centralized approach facilitates robust data management systems but raises significant concerns regarding data protection. In contrast, our FedSDAF framework addresses these privacy concerns by leveraging a federated learning paradigm, which allows for the exchange of knowledge without compromising client data.
Existing federated domain generalization (FedDG) methods, such as FedSR and FedADG, exhibit limitations, particularly in their reliance on local data without sufficient integration of domain-specific knowledge from diverse sources. While these methods have made strides in representation learning, they often fall short in achieving robust generalization across unseen domains. In contrast, our framework incorporates the utilization of source-domain perceptual features, which have proven instrumental in enhancing model adaptability to unseen target domains. This is particularly evident in our results, where the integration of domain-aware features has led to significant performance gains. Additionally, our bidirectional knowledge distillation mechanism facilitates effective knowledge sharing between clients, thereby enhancing model adaptability while preserving privacy.
4.5. Ablation Experiment
Table 4. TABLE VI INVESTIGATION ON THE CONTRIBUTION OF DIFFERENT COMPONENTS
BKD MHSA $\hat{A}_{di}$ $\hat{A}_{dw}$ PACS OfficeHome VLCS
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 98.52 87.63 91.38
$\times$ $\checkmark$ $\checkmark$ $\checkmark$ 97.46 85.81 86.19
$\checkmark$ $\times$ $\checkmark$ $\checkmark$ 97.13 85.42 87.20
$\checkmark$ $\times$ $\checkmark$ $\times$ 95.47 83.60 83.98
$\times$ $\times$ $\checkmark$ $\times$ 95.46 82.17 83.51
$\times$ $\times$ $\times$ $\times$ 93.29 78.70 82.69
To illustrate the significance of the various components utilized in the Federated Source Domain Awareness Framework (FedSDAF), we conducted an ablation study across three benchmark datasets. The results are presented in Table 4.
Optimal model performance was attained when all components were active, demonstrating the synergistic integration of features that collectively enhance the model’s generalization capacity. In contrast, disabling both the Bidirectional knowledge distillation (BKD) and the Multihead Self-Attention mechanism (MHSA) led to a significant decline in performance, highlighting the critical roles these components play in facilitating information exchange and feature extraction. The observed drop in accuracy indicates that, without these mechanisms, the model struggles to maintain the understanding necessary for effective generalization. This finding underscores the necessity of collaborative learning strategies, such as BKD, in optimizing model performance within federated contexts characterized by data isolation.
Further analysis revealed that disabling either adapter resulted in a decrease in model accuracy, emphasizing their integral roles within the architecture. Specifically, the reduction in performance upon removing the Domain-Invariant Adapter underscores its importance in extracting stable features crucial for robustness against domain shifts.The lowest performance was recorded when all components were disabled, confirming the complementary and synergistic nature of our framework. This substantial reduction in accuracy further emphasizes the necessity of each component’s functionality within the overall system.
4.6. Convergence Analysis
Figure 4. Convergence analysis on different clients in Domain benchmark
[…]
A woman and a girl in white dresses sit in an open car.
To validate the convergence properties of our FedSDAF framework, we conduct a comparative analysis against the baseline FedAVG (Adapter) method using the PACS dataset, which encompasses Art, Photo, Cartoon, and Sketch domains. As demonstrated in Figure 4, FedSDAF exhibits superior convergence characteristics in terms of both speed and stability across all domains, particularly during the initial communication rounds. FedSDAF demonstrates substantial early-stage convergence, achieving accuracy improvements of 5.8-7.2% over FedAVG within the first 10 communication rounds. This accelerated convergence can be attributed to the domain-aware adapter’s MHSA mechanism, which effectively prioritizes domain-specific feature extraction during local client updates. In contrast, FedAVG’s conventional adapter aggregation exhibits slower convergence, requiring additional rounds to reconcile heterogeneous domain representations due to its domain-agnostic nature. The convergence analysis extends to Round 40, where FedSDAF demonstrates stable convergence across all PACS domains with accuracy rates ranging from 93.6% to 97.1%, consistently outperforming FedAVG by margins of 9.4-13.8%. This enhanced convergence stability is primarily achieved through the bidirectional knowledge distillation mechanism, which facilitates systematic feature transfer between domains. Specifically, the framework’s ability to integrate diverse domain characteristics, such as texture-rich features from the Art domain with shape-centric patterns from the Sketch domain, contributes to its faster and more stable convergence trajectory. Unlike FedAVG’s static aggregation strategy, which often struggles to achieve stable convergence due to domain heterogeneity, FedSDAF’s iterative distillation process ensures consistent convergence behavior while effectively leveraging cross-domain complementarity.
4.7. Hyperparameter Analysis
Figure 5. Effects of $\alpha$.
We evaluate the robustness of FedSDAF through systematic analysis of the loss balancing coefficient $\alpha$ (Eq.7) across the entire PACS benchmark. FedSDAF consistently achieves accuracy exceeding 97% when $\alpha$ ranges from 0.3 to 5, with minimal variance (±0.8%). Notably, even with $\alpha$=0.1 (representing reduced distillation influence), the model maintains 96.2% accuracy, demonstrating the inherent stability of our dual-adapter architecture. Optimal performance is observed at $\alpha$=1, where domain-specific knowledge preservation and cross-client fusion reach an equilibrium. When $\alpha$ exceeds 5, a slight performance decline occurs (97.1% at $\alpha$=10), as the over-strengthened global distillation signal marginally inhibits local adaptation capabilities. These experimental results confirm FedSDAF’s robustness to hyperparameter selection, requiring only the straightforward choice of $\alpha$=1 to achieve peak performance without extensive tuning.
4.8. Visualization
Figure 6. Visualization of the target features extracted by ViT and FedSDAF.
[…]
A woman and a girl in white dresses sit in an open car.
We present several visualization results to provide an intuitive understanding of the advantages of FedSDAF. In Fig. 6, we apply the t-SNE algorithm (van der Maaten and Hinton, 2008) to visualize the features of target samples extracted by ViT-B/16, Domain Invariant Adapter, Domain Aware Adapter, and FedSDAF (KIA), respectively, with Sketch designated as the target domain on PACS. The t-SNE visualization reveals critical insights into FedSDAF’s feature discrimination capability within the challenging Sketch domain. While ViT-B/16 (Fig.6d) exhibits dispersed feature distribution with substantial inter-class overlap, both individual adapters demonstrate distinct adaptation strategies aligned with our design motivation: The Domain Invariant Adapter (Fig.6a) produces elliptical feature clusters with overlapping boundaries, reflecting its emphasis on cross-domain alignment at the expense of category discrimination. Conversely, the Domain Aware Adapter (Fig.6b) generates compact but partially misaligned clusters, indicative of its focus on capturing source-specific discriminative features while neglecting domain shift mitigation. FedSDAF (Fig.6c) fundamentally overcomes these limitations through hierarchical feature disentanglement. The isolated red cluster in the lower-left quadrant manifests the domain-aware component’s success in preserving source-discriminative patterns, while the uniform spacing between blue and purple clusters in the upper region verifies the domain-invariant module’s effectiveness in establishing cross-domain consistency. This dual mechanism achieves geometric synergy - the domain-aware adapter actively pushes class clusters apart by amplifying source-specific discriminators, while the domain-invariant adapter passively pulls features toward domain-agnostic decision boundaries. Such collaborative dynamics explain why FedSDAF surpasses both the base model and standalone adapters, as visualized by the sharpened cluster margins and reduced overlap density compared to other configurations. These visual patterns collectively validate our core motivation: Source-aware feature refinement requires simultaneous discriminative pattern preservation (through domain-aware adaptation) and cross-domain topology alignment (via invariant feature projection), which FedSDAF achieves through its dual-branch architecture.
5. Conclusion
In conclusion, the Federated Source Domain Awareness Framework (FedSDAF) represents a significant advancement in federated domain generalization by effectively leveraging source domain-aware features. This core contribution enhances model adaptation in environments characterized by data heterogeneity and privacy constraints. The dual adapter architecture, consisting of the Domain-Invariant Adapter and the Domain-Aware Adapter, is essential for maintaining stability while integrating domain-specific knowledge. This integration notably improves the model’s capacity to generalize across unknown target domains, addressing a key challenge faced by traditional federated learning methods.
However, certain limitations persist. The current study primarily focuses on specific domain datasets, which may not encompass the full variability encountered in real-world applications. Additionally, in this study, we transfer knowledge between the server and clients using logits of a public dataset, motivated by efficiency and privacy considerations. Although empirical evidence suggests that sharing logits of public datasets between the server and clients is more privacy-preserving than sharing model gradients or parameters (Fan et al., 2025; Liu et al., 2021c), there is no theoretical guarantee that this approach does not compromise the privacy of clients’ sensitive data. This issue warrants further investigation to ensure that privacy concerns are adequately addressed in future iterations of the framework.
Appendix A Research Methods
A.1. Part One:Recap: Generalization Adjustment Strategy
In order to mitigate the risk of extracted domain-invariant features being overly biased towards a specific source domain, we introduce a Generalization Adjustment (GA) strategy(Zhang et al., 2023c) during the aggregation phase. This section provides a formal description of the novel GA strategy designed for federated domain generalization. The GA mechanism incorporates a dynamic weight adjustment module on the server side, aimed at optimizing the global model’s generalization capability. The complete workflow encompasses four key phases.
Initially, in the Initialization Phase, The server initializes the global model parameters $\theta^{0}$ and sets uniform aggregation weights $a^{0}=(1/M,\ldots,1/M)$ for $M$ clients. Each client maintains a local copy of the model $\theta_{i}^{0}=\theta^{0}$ with domain-specific data $\widehat{D}_{i}$.
During the communication round $r$, the first step involves measuring the generalization gap. Each client computes the domain-specific generalization gap prior to local training, quantified by the equation:
(9) $\widehat{G}_{D}^{i}(\theta^{r})=\mathbb{E}^{\widehat{D}^{i}}(\theta^{r})-\mathbb{E}^{\widehat{D}^{i}}(\theta_{i}^{r-1})$
This equation captures the performance discrepancy between the current global model and the previous local optimum. Following this measurement, clients proceed with local model updates by performing E epochs of training using the algorithm to derive updated parameters
In the subsequent Server-side Weight Adjustment phase, the server undertakes two essential operations. The first is the momentum-based weight update, formulated as:
(10) $a_{i}^{r^{\prime}}=\left(\widehat{G}_{D}^{i}(\theta^{r})-\mu\right)\cdot\frac{d^{r}}{\max_{j}(\widehat{G}_{D}^{j}(\theta^{r})-\mu)}+a_{i}^{r-1}$
where $\mu$ denotes the average gaps and $d^{r}=(1-r/R)\cdot d$ implements a linear decay to control the update magnitude. Subsequently, the server conducts a generalization-aware aggregation, computed as:
(11) $\theta^{r+1}=\sum_{i=1}^{M}a_{i}^{r}\cdot\theta_{i}^{r^{\prime}}$
This reweighted aggregation approach supersedes the static weighting scheme employed by FedAvg.
Finally, in the Model Distribution phase, the refined global model $\theta^{r+1}$ is broadcast to all clients for subsequent rounds, thereby completing the closed-loop optimization cycle.
A.2. Part Two: Algorithm 1
As depicted in Algorithm 1 On the client side, each client initializes its local copy $\hat{A}_{di}^{k,r}$ corresponding to the received global domain-invariant adapter $A_{di}^{r}$. The client subsequently optimizes its local domain-aware adapter $A_{dw}^{k,r}$ and the local replica $\hat{A}_{di}^{k,r}$ through the bidirectional knowledge distillation mechanism. During this optimization, the client computes its training error $\hat{E}_{D_{k}}^{k,r}(A_{di}^{k,r})$ and, after each round, uploads the generalization gap $G_{k}^{r}$ (if $r>0$) along with the updated domain-invariant adapters $A_{di}^{k,r}$ back to the server. This systematic updating process enables the FedSDAF method to effectively leverage source domain knowledge, thereby enhancing the model’s generalization performance in unseen target domains.
Input: $K$: Number of clients
$\{D_{k}\}_{k=1}^{K}$: Client datasets
$R$: Communication rounds
Output: $\{A_{dw}^{k,R}\}_{k=1}^{K}$: Well-trained client models
$A_{di}^{R}$: Global domain-invariant adapter
1
2Server Execution:
3 Initialize $A_{di}^{0}$, $A_{dw}^{k,0}\leftarrow$ random initialization;
4 $a_{k}^{0}\leftarrow\frac{1}{M},\ \forall k\in\{1,\ldots,K\}$;
5
6for $r\leftarrow 0$ to $R-1$ do
7       if $r>0$ then
8             Collect $\{G_{k}^{r},A_{di}^{k,r}\}_{k=1}^{K}$;
9             $a_{k}^{r}\leftarrow\textsc{GA-Algorithm}(G_{k}^{r})$;
10             $A_{di}^{r+1}\leftarrow\sum_{k=1}^{K}a_{k}^{r}A_{di}^{k,r}$;
11
12       else
13             $A_{di}^{r+1}\leftarrow A_{di}^{0}$;
14
15       end if
16      Broadcast $A_{di}^{r+1}$;
17
18 end for
19
20Client $k$ Execution (Parallel):
21 for $r\leftarrow 0$ to $R-1$ do
22       Receive $A_{di}^{r}$;
23       $\hat{A}_{di}^{k,r}\leftarrow A_{di}^{r}$;
24       while not converged do
25             $\min_{A_{dw}^{k,r},\hat{A}_{di}^{k,r}}\mathcal{L}_{\text{distill}}(A_{dw}^{k,r},\hat{A}_{di}^{k,r})$;
26
27       end while
28      Compute $\hat{E}_{D_{k}}^{k,r}(A_{di}^{k,r})$;
29       if $r>0$ then
30             $G_{k}^{r}\leftarrow\hat{E}_{D_{k}}^{k,r}-\hat{E}_{D_{k}}^{k,r-1}$;
31
32       end if
33      Send $(G_{k}^{r},A_{di}^{k,r})$;
34
35 end for
Algorithm 1 Federated Domain Adaptation with Bidirectional Distillation (FedSDAF)